# APCTP-Pre2025-021

### v1

@@
We study systematically the total expansion experienced by a certain perturbation mode during single-field inflation, not resorting to explicit models of inflation or reheating. By assuming that during the reheating stage the equation of state $w_{\text{rh}}$ can be written as a function of $e$-folds, the unknown dynamics during reheating parametrized by $w_{\text{rh}}$ is confined within a time integral so that any dependence on the models of inflation and reheating is isolated from model-independent contributions. Especially, the dependence on the reheating dynamics via $w_{\text{rh}}$ and the reheating temperature $T_{\text{rh}}$ is dominating. We give two illustrative examples of $w_{\text{rh}}$ to discuss its impacts on the total expansion, which can be different as much as 10 even for the same reheating temperature, depending on the shape of $w_{\text{rh}}$.
@@ 1 Introduction
To derive inflation, viz. to make the expansion of the universe accelerating, we need a very unusual matter content that has a strong negative pressure. One such a matter content is the cosmological constant, which however does not lead to a graceful exit from inflation. Instead, we usually introduce a scalar field whose potential energy is dominant over the kinetic one so that it behaves effectively as a cosmological constant, leading to an accelerated, quasi-exponential expansion of the universe. This scalar field that dominates the energy density of the universe during inflation is called the inflaton [10, 11, 12].
Meanwhile, the standard paradigm for physical cosmology, the hot big bang model, describes successfully the evolution of the universe starting from big bang nucleosynthesis (see e.g. [13, 14]) once the initial conditions consistent with current observations – homogeneous and isotropic universe with ${\cal O}(10^{-5})$ fluctuations – are provided by inflation. This raises a very serious question. During inflation, the energy density of the universe is dominated by the inflaton. When the standard hot big bang evolution commences after inflation, relativistic matter contents such as photons occupy most of the energy density of the universe, along with tiny fraction of pressureless components like baryons and dark matter. This means the energy of the inflaton sector must be transferred to the usual matter contents with non-negative pressure, filling the universe with energetic particles. This process of energy transfer is called reheating (for reviews, see e.g. [15, 16, 17]).
Thus, without reheating, after inflation the universe never becomes hot and thermalized, not leading to the universe as observed. Reheating is an indispensable epoch that connects inflation and hot big bang. But as the universe, after reheating, is in thermal equilibrium, any model-dependent features that would have characterized the details of the reheating mechanism are swept away. This makes it very difficult to study the epoch of reheating. All we know is that the inflaton should decay and/or annihilate, and the final products are the standard model particles and dark matter. A standard, old approach is to introduce a decay width of the inflaton so that reheating proceeds perturbatively [18, 19, 20]. However, since reheating epoch is so elusive, any non-perturbative process may well take place, such as preheating [21, 22], rescattering [23, 24] and turbulence [25, 26, 27].
@@
In this article, we study systematically the contributions to the total expansion during inflation of a certain perturbation mode experiences. These contributions can be categorized into those dependent on and independent from the models of inflation and reheating dynamics. Only assuming that the equation of state during reheating can be written as a function of $e$-folds, we show that all the model dependence can be isolated from the model-independent contributions. This article is organized as follows. In Section 2, we consider the total expansion of a mode and separate individual contributions. By assuming that the equation of state during reheating is a function of $e$-folds, model-dependent contributions are sequestered from the model-independent ones. In Section 3 we study the two model-dependent inputs, the endpoint of inflation and the equation of state during reheating. We present a more accurate estimate on the endpoint given an inflation model, along with a simple, more phenomenological parametrization. We give explicit examples for the equation of state during reheating, and show the total expansion of a mode as a function of model parameters explicitly. Finally we conclude in Section 4.
@@ 2.1 Contributions to total expansion
    $N_{\text{rh}}$: This number measures the expansion of the universe starting from the end of inflation until reheating ends so that the universe becomes dominated by radiation. As $N_{\text{rh}}$ depends on the detail of the reheating mechanism that remains mostly elusive, we usually resort to an effective and/or phenomenological approach. One possible way is to make use of the equation of state during reheating, $w_{\text{rh}}$, and to assume that it is a constant, e.g. $w_{\text{rh}}=(n-2)/(n+2)$ for $V(\phi)\propto\phi^{n}$ when averaged over oscillation cycles [30]. This leads to the following relation between the energy densities at the end of inflation $\rho_{e}$ and at the end of reheating stage $\rho_{\text{rh}}$:
@@ 2.2 Sequestering model dependence
At the beginning of radiation domination, viz. at the end of reheating epoch, the equation of state is that of radiation: $w_{r}=1/3$. That is, at both ends of the reheating stage, the values of the equation of state are fixed, 1/3-1/3 at the beginning and $1/3$ at the end. Thus we expect that any realistic equation of state during reheating begins from 1/3-1/3, changes smoothly, and ends up at the value of 1/31/3.333Here, we assume that during the reheating state $w_{\text{rh}}$ does not become smaller than 1/3-1/3. If so, until $w_{\text{rh}}>-1/3$ again the universe undergoes yet another inflationary stage. Now we assume that $w_{\text{rh}}$ can be written as a function of $e$-folds, with $N=0$ at the beginning and $N=N_{\text{rh}}$ at the end of the reheating stage. Then $w_{\text{rh}}(N=0)=-1/3$ and $w_{\text{rh}}(N=N_{\text{rh}})=1/3$. From the continuity equation, we can solve for the energy density at the end of reheating, $\rho_{\text{rh}}$, with a general equation of state during reheating which is a function of the normalized $e$-folds $n\equiv N/N_{\text{rh}}$:
@@ Acknowledgements
We thank Kiwoon Choi, Suro Kim, Ryo Saito, Tomo Takahashi, Yuki Watanabe and Chul-Moon Yoo for helpful comments. This work is supported in part by Basic Science Research Program through the National Research Foundation of Korea (RS-2024-00336507). JG also acknowledges the Ewha Womans University Research Grant of 2025 (1-2025-0739-001-1). JG is grateful to the Asia Pacific Center for Theoretical Physics for hospitality during this work was under progress.

### v2

@@
We study systematically the total expansion experienced by a certain perturbation mode during single-field inflation, not resorting to explicit models of inflation or reheating. By assuming that during the reheating stage the equation of state $w_{\text{rh}}$ can be written as a function of $e$-folds, the unknown dynamics during reheating parametrized by $w_{\text{rh}}$ is confined within a time integral so that any dependence on the models of inflation and reheating is isolated from model-independent contributions. Especially, the dependence on the reheating dynamics via $w_{\text{rh}}$ and the reheating temperature $T_{\text{rh}}$ is dominating. We give two illustrative examples of $w_{\text{rh}}$ to discuss its impacts on the total expansion, which can be different as much as 10 even for the same reheating temperature, depending on the shape of $w_{\text{rh}}$. We also discuss the profile degeneracy of $w_{\text{rh}}$, and argue when the degeneracy is lifted.
@@ 1 Introduction
To drive inflation, viz. to make the expansion of the universe accelerating, we need a very unusual matter content that has a strong negative pressure. One such a matter content is the cosmological constant, which however does not lead to a graceful exit from inflation. Instead, we usually introduce a scalar field whose potential energy is dominant over the kinetic one so that it behaves effectively as a cosmological constant, leading to an accelerated, quasi-exponential expansion of the universe. This scalar field that dominates the energy density of the universe during inflation is called the inflaton [10, 11, 12].
Meanwhile, the standard paradigm for physical cosmology, the hot big bang model, describes successfully the evolution of the universe starting from big bang nucleosynthesis (see e.g. [13, 14]) once the initial conditions consistent with current observations – homogeneous and isotropic universe with ${\cal O}(10^{-5})$ fluctuations – are provided by inflation. This raises a very serious question: How does the universe become so hot after “cold” inflation? During inflation, the energy density of the universe is dominated by the inflaton. When the standard hot big bang evolution commences after inflation, relativistic matter contents such as photons occupy most of the energy density of the universe, along with tiny fraction of pressureless components like baryons and dark matter. This means the energy of the inflaton sector must be transferred to the usual matter contents with non-negative pressure, filling the universe with energetic particles. This process of energy transfer is called reheating (for reviews, see e.g. [15, 16, 17]).
Thus, without reheating, after inflation the universe never becomes hot and thermalized [^3] Precisely speaking, this statement is true only if radiation is produced after inflation. If radiation exists during inflation in e.g. warm inflation [18] the post-inflationary evolution including reheating stage could be different from the standard, “cold” inflation scenario (see e.g. [19]). , not leading to the universe as observed. Reheating is an indispensable epoch that connects inflation and hot big bang. But as the universe, after reheating, is in thermal equilibrium, any model-dependent features that would have characterized the details of the reheating mechanism are swept away. This makes it very difficult to study the epoch of reheating. All we know is that the inflaton should decay and/or annihilate, and the final products are the standard model particles and dark matter. A standard, old approach is to introduce a decay width of the inflaton so that reheating proceeds perturbatively [20, 21, 22]. However, since reheating epoch is so elusive, any non-perturbative process may well take place, such as preheating [23, 24], rescattering [25, 26] and turbulence [27, 28, 29].
@@
In this article, we study systematically the contributions to the total expansion during inflation of a certain perturbation mode experiences. These contributions can be categorized into those dependent on and independent from the models of inflation and reheating dynamics. Only assuming that the equation of state during reheating can be written as a function of $e$-folds, we show that all the model dependence can be isolated from the model-independent contributions. This article is organized as follows. In Section 2, we consider the total expansion of a mode and separate individual contributions. By assuming that the equation of state during reheating is a function of $e$-folds, model-dependent contributions are sequestered from the model-independent ones. In Section 3 we study the two model-dependent inputs, the endpoint of inflation and the equation of state during reheating. We present a more accurate estimate on the endpoint given an inflation model, along with a simple, more phenomenological parametrization. We give explicit examples for the equation of state during reheating, and show the total expansion of a mode as a function of model parameters explicitly. In Section 4, we argue that there exist degeneracies in different profiles of $w_{\text{rh}}$, the equation of state during reheating, as long as the average value of $w_{\text{rh}}$ is identical, and illustrate when and how the degeneracy is lifted and what information we can gain regarding reheating. Finally we conclude in Section 5.
@@ 2.1 Contributions to total expansion
    $N_{\text{rh}}$: This number measures the expansion of the universe starting from the end of inflation until reheating ends so that the universe becomes dominated by radiation. As $N_{\text{rh}}$ depends on the detail of the reheating mechanism that remains mostly elusive, we usually resort to an effective and/or phenomenological approach. One possible way is to make use of $w_{\text{rh}}$ and to assume that it is a constant, e.g. $w_{\text{rh}}=(n-2)/(n+2)$ for $V(\phi)\propto\phi^{n}$ when averaged over oscillation cycles [32]. This leads to the following relation between the energy densities at the end of inflation $\rho_{e}$ and at the end of reheating stage $\rho_{\text{rh}}$:
@@ 2.2 Sequestering model dependence
At the beginning of radiation domination, viz. at the end of reheating epoch, the equation of state is that of radiation: $w_{r}=1/3$. That is, at both ends of the reheating stage, the values of the equation of state are fixed, 1/3-1/3 at the beginning and $1/3$ at the end. Thus we expect that any realistic equation of state during reheating begins from 1/3-1/3, changes smoothly, and ends up at the value of 1/31/3.444Here, we assume that during the reheating state $w_{\text{rh}}$ does not become smaller than 1/3-1/3. If so, until $w_{\text{rh}}>-1/3$ again the universe undergoes yet another inflationary stage. Now we assume that $w_{\text{rh}}$ can be written as a function of $e$-folds, with $N=0$ at the beginning and $N=N_{\text{rh}}$ at the end of the reheating stage. Then $w_{\text{rh}}(N=0)=-1/3$ and $w_{\text{rh}}(N=N_{\text{rh}})=1/3$. From the continuity equation, we can solve for the energy density at the end of reheating, $\rho_{\text{rh}}$, with a general equation of state during reheating which is a function of the normalized $e$-folds $n\equiv N/N_{\text{rh}}$:
@@ 3.2 Equation of state during reheating
4 Degeneracy with the effective equation of state
Until now, we have considered the separation of the model-dependent and weakly independent contribution to the total expansion. As can be read from (2.2), the detail during the reheating stage is confined within $\cal W$}, with $\rho_{e}$ and $T_{\text{rh}}$ being the fixed boundaries at the beginning and end of the reheating stage respectively. Thus, if we define the average equation of state $w_{\text{eff}}$ as (see e.g. [39])
$w_{\text{eff}}\equiv\frac{1}{N_{\text{rh}}}\int_{0}^{N_{\text{rh}}}w_{\text{rh}}(N)dN=\int_{0}^{1}w_{\text{rh}}(n)dn={\cal W}\,$, (32)
there is no difference at all from just using $w_{\text{eff}}$ instead of $\cal W$}. That is, the information on the reheating stage contained in $N_{k}$ is only the average value of $w_{\text{rh}}$. More precisely speaking, given a model of inflation specified by $\rho_{e}$, assuming a standard thermal history with constant relativistic degrees of freedom $g_{*}$, the reheating contribution to $N_{k}$ depends only on the averaged equation of state $w_{\text{eff}}$ and the reheating temperature $T_{\text{rh}}$. In this case, different reheating profiles $w_{\text{rh}}(N)$ that share the same $w_{\text{eff}}$ are exactly degenerate in their prediction for $N_{k}$. Our parameter $\alpha$ that specifies the evolution of $w_{\text{rh}}$ during reheating therefore does not introduce an additional degree of freedom for itself in this case. However, this does not mean that using the full profile of $w_{\text{rh}}$ is useless. Rather, what we have seen above indicates that if we choose to compute $N_{k}$ alone, the information on the details of $w_{\text{rh}}$ are collapsed into the average value $w_{\text{eff}}$: $N_{k}$ alone is blind to the shape of reheating.
4.1 When the degeneracy is broken
Therefore, the shape of $w_{\text{rh}}$ becomes relevant when our assumptions are violated: During reheating the universe is described by a single barotropic fluid with a constant $g_{*}$, and we are only interested in $N_{k}$. Thus, there are largely two different ways to break the profile degeneracy.
  1. 1.
    As can be read from the background solution (10), all background observables such as $N_{k}$ are functionals of $w_{\text{rh}}(N)$ only through the single number $w_{\text{eff}}$. This is because the Friedmann-Robertson-Walker (FRW) background acts as a “projection operator” that maps the infinite-dimensional function space of $w_{\text{rh}}(N)$ onto a one-dimensional parameter space of $w_{\text{eff}}$. Thus, The profile of $w_{\text{rh}}(N)$ becomes relevant when the observable probes the operators beyond the background stress-energy tensor. More precisely speaking, the background evolution depends only on $T_{\mu\nu}$ evaluated on the homogeneous solution. However, any observables that probe local time dependence during reheating, such as horizon reentry of sub-CMB modes, perturbation evolution, or particle production depend on quantities like $H(N)$, $dH/dN$, and so on, which again depend on the local value of $w_{\text{rh}}(N)$, not its integral. For example, the gravitational wave spectrum $\Omega_{\text{GW}}h^{2}$ is sensitive to the profile of $w_{\text{rh}}$ (see e.g. [40]).
    We can more formally state the above discussions using the terms of effective field theory (EFT). The background and perturbation operators can be written in the effective action as
    $S_{\text{eff}}=\int d^{4}x\sqrt{-g}\Big[{\cal L}_{\text{bg}}(w_{\text{eff}})+{\cal L}_{\text{pert}}(w_{\text{rh}}(N),c_{s}(N),\cdots)+\cdots\Big]\,.$ (33)
    As the background sector depends only on $w_{\text{eff}}$, the functional profile of $w_{\text{rh}}(N)$ is invisible to the EFT truncated to homogeneous operators. We can access to the profile dependence and hence break degeneracy of $w_{\text{eff}}$ only when the perturbation operators are retained that resolve local time dependence during reheating.
  2. 2.
    Incorporating other observables, such as $\Omega_{\text{GW}}h^{2}$, along with $N_{k}$ is clearly one way of breaking the profile degeneracy of $w_{\text{rh}}$. There is yet another. We previously have assumed more than just FRW and energy conservation. We have also assumed that the post-reheating radiation bath behaves adiabatically with effectively constant relativistic degrees of freedom $g_{*}$. Thus, for example, once $g_{*}$ varies during reheating, the background reduction to $w_{\text{eff}}$ is no longer exact.
    In writing (10), we have only assumed barotropicity $p=w_{\text{rh}}\rho$ during reheating, and no assumption about $g_{*}$ is made. But when we translate it into a relation involving temperature, we use the energy density for radiation
    $\rho_{r}=\frac{\pi^{2}}{30}g_{*}(T)T^{4}\,$, (34)
    and entropy conservation
    $g_{*}(T)a^{3}T^{3}=\text{constant}\,.$ (35)
    Hence, if $g_{*}$ varies significantly during reheating, then
    $d\log{T}=-dN-\frac{1}{3}d\log g_{*}\,$, (36)
    so that the mapping between $a$, $T$ and $\rho$ picks up explicit dependence on the local evolution of $g_{*}$, which cannot be absorbed into $w_{\text{eff}}$.
    The varying $g_{*}$ means that additional degrees of freedom become dynamical, or that phase transition occurs, or that entropy is produced non-adiabatically and redistributed. So even at homogeneous level, the effective stress-energy tensor is no longer characterized by a single parameter $w_{\text{eff}}$. Thus, now the background action in (33) depends also on $g_{*}(N)$, and different reheating histories with identical $w_{\text{eff}}$ can yield different final temperatures and different mappings between $k$ and $N_{k}$. In this sense, “weak model-independence” means setting up the boundaries up to which point the EFT description of reheating, parametrized by $w_{\text{eff}}$, is valid: It is so within a) a single perfect fluid, b) adiabatic evolution, and c) fixed particle contents.
4.2 Shape sensitivity from varying relativistic degrees of freedom
Before we proceed with a physically motivated simple study, we first consider a toy MCMC case for illustration how $N_{k}$ depends on the shape parameter $\alpha$. With a fiducial value $N_{k}^{\text{fid}}$, we add to $N_{k}$ the part $\delta{N}_{k}$ that depends on $\alpha$:
$N_{k}=N_{k}^{\text{fid}}+\delta{N}_{k}(\alpha)\,.$ (37)
As a minimal choice, let us consider a linear toy dependence:
$\delta{N}_{k}(\alpha)=A\alpha\,$, (38)
where $A$ is some number. We choose the likelihood of $N_{k}$ to be Gaussian with the variance $\sigma$, centered on the fiducial value $N_{k}^{\text{fid}}$. In Figure 3, we show the posterior distribution of the shape parameter $\alpha$, chosen randomly between 1-1 and 1, obtained from the MCMC analysis with 50,000 samples incorporating (38). The strong Gaussian-like constraint centered near $\alpha=0$ with exponential suppression beyond $|\alpha|\sim 0.25$ clearly demonstrates that the reheating profile shape is non-degenerate and can be constrained with precise measurements, addressing the limitation of the standard parameterization using only $w_{\text{eff}}$.
Figure 3: Posterior distribution of $\alpha$ obtained from the MCMC analysis for (38). We have set $N_{k}^{\text{fid}}=55$, $A=3.0$ and $\sigma=0.2$. We have taken 50,000 samples with $1<\alpha<1$.
Now, we consider an explicit toy model where $g_{*}$ varies as a step:
$g_{*}(T)=\left\{\begin{array}[]{ll}g_{h}&(T>T_{c})\vskip 3.00003pt\\ g_{l}&(T<T_{c})\end{array}\right$.\,, (39)
with $g_{h}>g_{l}$. We set up two different profiles for $w_{\text{rh}}(n)$ as
$w_{a}(n)$ $=\left\{\begin{array}[]{ll}w_{s}&(0<n<\alpha)\vskip 3.00003pt\\ w_{r}&(\alpha<n<1)\end{array}\right$.\,, (42)
$w_{b}(n)$ $=\left\{\begin{array}[]{ll}w_{r}&(0<n<1-\alpha)\vskip 3.00003pt\\ w_{s}&(1-\alpha<n<1)\end{array}\right$.\,, (45)
with $w_{s}>w_{r}$. That is, we give different ordering of stiff and soft phases: Stiff phase comes early ($w_{a}$) or late ($w_{b}$). In both cases, the average equation of state $w_{\text{eff}}$ is the same:
$w_{\text{eff}}=(w_{s}-w_{r})\alpha+w_{r}\,$, (46)
so that the $\alpha$-dependence of $N_{k}$ can be, at least approximately, modeled by (38). Because of entropy conservation (35), when $g_{*}$ drops from $g_{h}$ to $g_{l}$, temperature does not scale simply as $T\propto a^{-1}$, but there is an extra jump $T\propto g_{*}^{-1/3}a^{-1}$. This makes the background mapping ordering-sensitive, not just average-sensitive.
First we begin with the energy density evolution (10). Since $w_{\text{eff}}={\cal W}$ is identical for both profiles, the energy density evolves identically so we have the same $\rho_{\text{rh}}$. Up to here, it is still degenerate. Now suppose that the temperature crosses $T_{c}$ at some $e$-folds $N_{c}$ during reheating. From (34), at the crossing we have
$\rho(N_{c})=\frac{\pi^{2}}{30}g_{h}T_{c}^{4}\,.$ (47)
Because $\rho$ depends on the integrated history up to that time, the value of $N_{c}$ depends on how much stiff versus soft expansion happened before the crossing as during the stiff phase $\rho$ decreases faster. So, If the stiff phase happens early ($w_{a}$), the energy density drops faster before hitting $T_{c}$, and $N_{c}$ occurs earlier. If, conversely, the stiff phase comes late ($w_{b}$), the crossing happens later. Thus, even though $w_{\text{eff}}$ is degenerate, the moment of crossing $N_{c}$ becomes different depending on the profile of $w_{\text{rh}}$. This gives rise to different pair $(a_{\text{rh}},T_{\text{rh}})$ for different profiles.
@@ 5 Conclusions
Moreover, we have clarified on very general ground when the degeneracy in the shape profile of $w_{\text{rh}}$ matters. There are two ways of breaking degeneracy. One is to incorporate another perturbation observable, and the other is to vary the relativistic degrees of freedom $g_{*}$. We have shown how degeneracy can be broken by presenting an explicit example of varying $g_{*}$.
@@ Acknowledgements
We thank Kiwoon Choi, Suro Kim, Ryo Saito, Tomo Takahashi, Yuki Watanabe and Chul-Moon Yoo for helpful comments. We are also grateful to the anonymous editor for her/his constructive feedback that improved the contents of this article. This work is supported in part by Basic Science Research Program through the National Research Foundation of Korea (RS-2024-00336507). JG also acknowledges the Ewha Womans University Research Grant of 2025 (1-2025-0739-001-1). JG is grateful to the Asia Pacific Center for Theoretical Physics for hospitality during this work was under progress.